GlueX Start Counter Meeting, May 29, 2014

From GlueXWiki
Jump to: navigation, search

GlueX Start Counter Meeting
Thursday, May 29, 2014
10:00 am EDT
JLab: CEBAF Center F326


  1. Announcements
  2. Review minutes from previous meeting
  3. Detector Construction Status (Eric)
  4. Support Structure Construction Status (Tim/Chuck)
  5. Schedule Review (all)
  6. Software Status (all)

Communication Information

We will use BlueJeans for this meeting.

  • To join or start the meeting, go here.
  • Connecting directly from a room system? (E. g. Polycom H.323 device.)
    1. Dial: or
    2. Enter Meeting ID: 556286544 -or- use the pairing code
  • Just want to dial in?
    1. Dial-in: +1 408 740 7256 or +1 888 240 2560 (US or Canada only) (other phone numbers)
    2. Enter Meeting ID: 556286544


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2014-2Q on the JLab CUE. This directory is accessible from the web at .



  • FIU: Werner Boeglin, Eric Pooser
  • JLab: Fernando Barbosa, Eugene Chudakov, Chuck Hutton, Mark Ito (chair), Nick Sandoval, Beni Zihlmann

Nick's Trip to FIU

  • We set a target date of the week after the Fourth of July holiday.
  • Werner has all of the read-out electronics he needs at FIU. The MPOD does need to be shipped to support testing.
  • We decided to ship all of the cables that connect ST2 to ST3 to FIU.
  • We will ship 10 sets of cables (for 10 counters) for testing. These are the full length cables that will be used in the Hall. Fernando points out that this will give more realistic test results.
  • Nick will bring a hand-held 4-channel thermister read-out device with him. Other multi-meters are available at FIU for use after he leaves.

Construction Status

Eric gave the report.

  • The 50 scintillators are on their way to FIU from McNeal. They should arrive later today.
  • Eric has looked at the measurement data from McNeal. They look nominal.
  • Ibrahim at McNeal told Eric that there was no need to remove the protective adhesive machining.
  • The assembly jig is ready to go.

Support Structure Construction Status

Chuck gave the report.

  • The aluminum support rings are in hand. A test fit of the electronics on one of the rings was successful.
  • The final machined Rohacell is due in tomorrow.
  • The carbon fiber inner support is ready to go.
  • The support assembly will all be glued together next week.

Schedule Review

We thought that the readiness review should be scheduled for the end of June, before Nick's trip. By then FIU will have had an initial look at any assembly issues.

Testing Individual SiPMs

We discussed techniques for ensuring that all four SiPMs in a read-out channel are contributing to the signal. All sets of four share a single bias voltage and are summed into a single pre-amp. The tests with the laser used a diffuser that illuminated all four at the same time. There is also concern that during the various assembly stages, connections on the board my get broken. When the detector is completely assembled there access to the SiPMs one at a time is blocked by the scintillator counters themselves. We decided on the following steps.

  • Nick will devise a method for doing a laser test of the SiPMs one at a time. This tells us that all four SiPMs are contributing, at least at the pre-assembly stage.
  • We will record record current for each channel. If these shift in the near future, then that will be an indication that a SiPM has become uncoupled. Over the long-term it may to hard to maintain this as a useful technique due to changes in the SiPMs themselves, due to radiation for example.

Mark pointed out that we made some decisions early on to save money and simplify construction that preclude simple checking of the performance of individual SiPMs and that this was recognized at the time. Nothing we know now invalidates those conclusions.

Bias Voltage Control

Fernando asked about the the control system for the bias voltages. Mark will check with Hovanes Egiyan on the status.

Light Tighting Issues

Eric and Chuck have discussed the issue of keeping the detector light tight from the upstream end. We have to make a definite plan on how to do this and still allow for air flow to cool the electronics.